# StoryTeller: Improving Long Video Description through Global Audio-Visual Character Identification

### v2

@@ 5.1 Baselines
  • Closed-Source LVLMs: We compare our method with Gemini-1.5-pro [38] and GPT-4o [32]. For both models, we perform prompt engineering to generate detailed video descriptions that incorporate both fine-grained details and high-level plot information. The prompt templates are provided in Appendix F.
  • Open-Source LVLMs: We also compare our method with open-source LVLMs that demonstrate strong video understanding capabilities, including VILA1.5-8B [27], LLaVA-OneVision-7B [23], Qwen2-VL-7B [34], and InternVL2-8B [7, 8]. To ensure a fair comparison, we fine-tune each model using MovieStory101 training set.
Model StoryQA Accuracy
Character Action Plot Total
Gemini-1.5-pro 0.578 0.501 0.534 0.544
GPT-4o 0.517 0.479 0.528 0.507
VILA1.5-8B 0.561 0.459 0.540 0.524
LLaVA-OneVision-7B 0.557 0.454 0.540 0.520
Qwen2-VL-7B 0.549 0.468 0.549 0.523
InternVL2-8B 0.535 0.448 0.506 0.501
StoryTeller 0.676 0.583 0.644 0.639
Table 4: The results of video descriptions generated by StoryTeller and baselines evluated on StoryQA.
Human: Principal said seriously, “Come out for a moment.” Teacher Wang said anxiously: “Head to self-study.” Just as Teacher Wang left, Xia Luo picked up the textbook and approached Yuan Hua. “Move to the back,” he instructed. Yuan Hua replied irritably: “Did Teacher Wang agree to this?” Xia Luo said arrogantly, “If he doesn’t, I’ll just flatter him again.” Yuan Hua glared at Xia Luo, while Qiu Ya, looking worried, asked, “What should I do, Yuan Hua?” With a smile, Yuan Hua said, “Don’t worry, I’ll be back.” Yuan Hua walked to the last row, and Xia Luo happily took the seat next to Qiu Ya. Ma Dongmei also moved to the seat behind Qiu Ya. Ma Dongmei looked at Meng Te and said, “You go to the back.” Meng Te said fearfully, “Ma Dongmei, You can’t tear us apart just to pursue your own happiness.” Ma Dongmei broke Meng Te’s pencil on the spot…
StoryTeller: Principal looked at Teacher Wang and said, “Come out for a moment.”. Principal left the classroom, Teacher Wang turned around and said, “Head to self-study,” then Teacher Wang also left the classroom. Xia Luo rolled his eyes, grabbed his books, and walked towards the back of the classroom. He patted Yuan Hua’s shoulder and said, “Move to the back.” Yuan Hua asked: “Did Teacher Wang agree to this?” Xia Luo continued: “If he doesn’t, I’ll just flatter him again.” Yuan Hua, who was sitting behind, overheard Xia Luo’s words, then turned and whispered to Qiu Ya, who was sitting in front of him. Qiu Ya asked, “What should I do, Yuan Hua?” Yuan Hua looked at Qiu Ya and said, “Don’t worry, I’ll be back.” Yuan Hua left after speaking, Xia Luo sat down and looked at Qiu Ya. Ma Dongmei said, “You go to the back.” Meng Te turned to Ma Dongmei and said, “Ma Dongmei, You can’t tear us apart just to pursue your own happiness.” Ma Dongmei snapped the pen in two and threw it on the ground
Gemini-1.5-pro: Principal calls Xia Luo out of the classroom. Teacher Wang asks Ma Dongmei to go to self-study. Xia Luo walks to the back of the classroom and stands next to Yuan Hua. Xia Luo asks Yuan Hua to change seats with him and threatens to beat him if he doesn’t agree. Yuan Hua looks back at his girlfriend, Qiu Ya, with a concerned expression. Xia Luo changes seats with Yuan Hua and sits next to Qiu Ya. Yuan Hua worries about Qiu Ya and asks her what to do. Xia Luo assures Yuan Hua that he will come back eventually. Yuan Hua smiles reluctantly and walks to the back of the classroom. Xia Luo smiles at Qiu Ya. She looks back at him with an annoyed expression. Ma Dongmei walks towards Qiu Ya and scolds her for breaking them up. She throws a pen in anger and Yuan Hua is upset
Table 5: Compare StoryTeller and Gemini-1.5-pro on a video example featuring multiple characters. Text in green highlights accurate details that are missed by the other model, while text in red indicates hallucinated descriptions. Note that the displayed video frames are manually selected key frames for illustrative purposes and are not equivalent to the input frames used by the models.
@@
Comparison StoryTeller wins Same Baseline wins Advantage
StoryTeller vs Gemini-1.5-pro 40.89% 33.78% 25.33% +15.56%
StoryTeller vs VILA1.5-8B 56.11% 30.03% 13.86% +42.25%
Table 6: Human side-by-side evaluation results on the MovieStory101 test set. The advantage rate is calculated as the difference between the win rate and the loss rate.
@@ 5.2 Experimental Results
Gemini-1.5-pro[38] - 0.544 - 0.566 4.0% 0.574 1.4%
GPT-4o[32] 0.507 0.519 2.4% 0.564 8.7% 0.573 1.6%
VILA1.5-8B[27] 0.524 0.566 8.0% 0.6 6.0% 0.605 0.8%
LLaVA-OneVision-7B[23] 0.52 0.561 7.9% 0.593 5.7% 0.598 0.8%
Qwen2-VL-7B[34] 0.523 0.555 6.1% 0.597 7.6% 0.601 0.7%
InternVL2-8B[7, 8] 0.501 0.541 8.0% 0.578 6.8% 0.582 0.7%
@@
Table 7: Evaluation of StoryTeller as a general framework for improving video descriptions across vairous LVLMs. The performance on StoryQA when key components are incrementally added: video segmentation module, integration of the audio-visual character identification (with ‘Others’ label for unlisted characters), and the use of descriptive names for unlisted characters. The baseline approach involves segmenting videos into 10-second clips and generating descriptions without audio-visual character identification.
@@
The system design of StoryTeller is a general framework applicable to various LVLMs to enhance their long video descriptions. Table 7 presents the additive contributions of each component for various LVLMs. First, replacing the direct division into 10-second clips with our video segmentation module consistently improves description accuracy across all models, resulting in an average relative increase of 6.3% in accuracy on the StoryQA. Second, utilizing the audio-visual identification results generated by our model as input for the LVLMs, we observe an average relative accuracy improvement of 6.4%. This demonstrates that our model’s audio-visual identification results is effective for enhancing video descriptions across various LVLMs. Lastly, we observe that replacing the label ‘Others’ with descriptive names generated by our model further improves video descriptions, consistently benefiting all models. Notably, under our framework, Gemini-1.5-pro and GPT-4o achieve relative improvements of 5.5% and 13.0% in accuracy on StoryQA, respectively.
@@ 6 Conclusion
In this paper, we figure out audio-visual character identification as a crucial step for generating detailed and consistent descriptions of long videos. Based on this insight, we propose StoryTeller, a system for producing dense, plot-level descriptions, composed of three modules: video segmentation, audio-visual character identification, and description generation. We also introduce a new dataset, MovieStory101, with 5,982 three-minute movie clips from 187 movies. To effectively and efficiently evaluate long video descriptions, we present StoryQA, a large-scale multiple-choice question-answering dataset aligned with the MovieStory101 test set. We assess descriptions by inputting them into GPT-4 to answer these questions, using accuracy as an automatic evaluation metric. Experiments demonstrate that StoryTeller outperforms all baselines, surpassing Gemini-1.5-pro by 9.5% in accuracy on StoryQA and achieving a +15.56% advantage in human side-by-side evaluations. Additionally, we verify that incorporating audio-visual character identification significantly enhances video descriptions for all LVLMs, with Gemini-1.5-pro and GPT-4o achieving accuracy improvements of 5.5% and 13.0%, respectively.
Appendix A Annotation Guidelines for Movie Description
@@ Appendix C Long Video Segmentation Method
  1. 1.
    Use the PySceneDetect[2] to identify a series of cut points in the video.
  2. 2.
@@ Appendix D Global ID Generation
Audio Embedding. We fine-tune the ERes2NetV2 model [5] to obtain improved audio embeddings for our scenario. The training data is based on the MovieStory101 training set, from which we construct a series of pairwise samples: positive pairs consist of two audio segments from the same person, while negative pairs consist of segments from different individuals. In total, we collect 246,834 data pairs, with 5% randomly sampled as the development set, and the rest used for training. The ERes2NetV2 model is fine-tuned using CosineEmbeddingLoss as the loss function. We set the learning rate to 0.005 and the batch size to 32.
@@
Additionally, our audio embedding model outperforms all open-source models under this condition, as Table 10 shown.
@@
CAM++[43] 0.22 0.36
Ecapa-tdnn[10] 0.23 0.36
ERes2NetV2[5] 0.25 0.40
@@ Appendix E Training Details
Pre-training the audio module. In the pre-training stage, we use the same pre-training dataset of SALMONN [40], consisting of both 960-hour LibriSpeech training set [33] and 1000-hour GigaSpeech M-set [3] for speech recognition, as well as 2800-hour WavCaps [29], AudioCaps [21] and Clotho [11] for audio captioning. In this phase, we only pre-train the window-level Q-former, while freezing the audio encoders and LLM. For the window-level Q-former, we use only one trainable query, and set $L=17$, which corresponds to approximately 0.33 seconds per window. This configuration outputs 88 textual tokens from the Q-former for a 30-second audio segment. The model was trained on 16 H800 GPUs for 50,000 steps using a batch size of 6.
Fine-tuning on audio and visual tasks. To enhance the model’s ability in audio and visual character identification, we collect training data for the following tasks: audio diarization, and character recognition, identification, and tracking. For audio diarization, we sampled data from AMI [28], AISHELL-4 [12], AliMeeting [47, 48], CallHome [1], DIHARD [39], VoxConverse [9], and LibriCSS [6], creating a training dataset consisting of 290,088 minute-level utterances. Additionally, using MovieNet [18], we construct training data for three tasks: (1) character recognition, which involves recognizing characters from frames; (2) character identification, where given a cast list, the model identifies the characters in the frames; and (3) character tracking, where the model tracks a character’s position across all frames based on a given face. We collect 33,576 data for these three tasks.

### v3

@@ 5.1 Baselines
  • Closed-Source LVLMs: We compare our method with Gemini-1.5-pro (Reid et al. 2024) and GPT-4o (OpenAI 2024). For both models, we perform prompt engineering to generate detailed video descriptions that incorporate both fine-grained details and high-level plot information. The prompt templates are provided in Appendix F.
  • Open-Source LVLMs: We also compare our method with open-source LVLMs that demonstrate strong video understanding capabilities, including VILA1.5-8B (Lin et al. 2024), LLaVA-OneVision-7B (Li et al. 2024), Qwen2-VL-7B (Peng et al. 2024), and InternVL2-8B (Chen et al. 2023, 2024b), all fine-tuned on the MovieStory101 training set for fairness.
@@ 5.2 Experimental Results
Baseline Win Same Loss Advantage
Gemini-1.5-pro 40.89% 33.78% 25.33% +15.56%
VILA1.5-8B 56.11% 30.03% 13.86% +42.25%
Table 5: Human side-by-side evaluation results on the MovieStory101 test set (StoryTeller vs. baseline). Win, Same, and Loss refer to StoryTeller outcomes. Advantage is calculated as WinLoss.
@@
Gemini-1.5-pro(Reid et al. 2024) - 0.544 - 0.566 4.0% 0.574 1.4%
GPT-4o(OpenAI 2024) 0.507 0.519 2.4% 0.564 8.7% 0.573 1.6%
VILA1.5-8B(Lin et al. 2024) 0.524 0.566 8.0% 0.6 6.0% 0.605 0.8%
LLaVA-OneVision-7B(Li et al. 2024) 0.52 0.561 7.9% 0.593 5.7% 0.598 0.8%
Qwen2-VL-7B(Peng et al. 2024) 0.523 0.555 6.1% 0.597 7.6% 0.601 0.7%
InternVL2-8B(Chen et al. 2023, 2024b) 0.501 0.541 8.0% 0.578 6.8% 0.582 0.7%
@@
Table 6: Evaluation of StoryTeller as a general framework for improving video descriptions across vairous LVLMs. The performance on StoryQA when key components are incrementally added: video segmentation module, integration of the audio-visual character identification (with ‘Others’ label for unlisted characters), and the use of descriptive names. The baseline approach involves segmenting videos into 10-second clips and generating descriptions without audio-visual character identification.
@@
Besides StoryQA, we report results from traditional evaluation metrics, including CIDEr (Vedantam, Lawrence Zitnick, and Parikh 2015), R@1/5 (Han et al. 2023a) and CRITIC (Han et al. 2024), as shown in Figure 4. We should note that traditional metrics (e.g., BLEU, ROUGE, CIDEr) rely on n-gram overlap, often penalizing accurate descriptions with different wording. Embedding-based metrics like R@1/5 improve semantic evaluation but still struggle with long, detailed content. This limitation is evident in Table 4, where Gemini-1.5-pro scores lower than VILA1.5-8B despite outperforming it in human side-by-side evaluations. At last, as modern video description models are largely LLM-based and generally produce fluent outputs, we prioritize evaluating content accuracy. StoryQA serves as an effective and scalable automated evaluation approach.
The system design of StoryTeller is a general framework applicable to various LVLMs to enhance their long video descriptions. Table 6 presents the additive contributions of each component for various LVLMs. First, replacing the direct division into 10-second clips with our video segmentation module consistently improves description accuracy across all models, resulting in an average relative increase of 6.3% in accuracy on the StoryQA. Second, utilizing the audio-visual identification results generated by our model as input for the LVLMs, we observe an average relative accuracy improvement of 6.4%. This demonstrates that our model’s audio-visual identification results is effective for enhancing video descriptions across various LVLMs. Lastly, we observe that replacing ‘Others’ with descriptive names generated by our model further improves video descriptions, consistently benefiting all models. Notably, under our framework, Gemini-1.5-pro and GPT-4o achieve relative improvements of 5.5% and 13.0% in accuracy, respectively.
@@ 6 Conclusion
In this paper, we introduce StoryTeller, a system for producing detailed, plot-level descriptions, composed of three modules: video segmentation, audio-visual character identification, and description generation. We introduce a new dataset, MovieStory101, with 5,982 three-minute movie clips from 187 movies. To effectively and efficiently evaluate long video descriptions, we present StoryQA, a large-scale multiple-choice question-answering dataset aligned with the MovieStory101 test set. We assess descriptions by inputting them into GPT-4 to answer these questions, using accuracy as an automatic evaluation metric. Experiments demonstrate that StoryTeller outperforms all baselines, surpassing Gemini-1.5-pro by 9.5% in accuracy on StoryQA and achieving a +15.56% advantage in human side-by-side evaluations.
Appendix A Annotation Details
A.1 Annotation Guidelines for Movie Description
@@
A.2 Annotator Information
All annotators are employed by a commercial data annotation company. We sign a contract with the company and pay the company for the annotation work at a market price. The annotators are all college graduates with strong English proficiency. A total of 15 annotators participated in the annotation work.
@@ Appendix C Long Video Segmentation Method
  • Use the PySceneDetect(Castellano 2014) to identify a series of cut points in the video.
@@ Appendix D Global ID Generation
Audio Embedding. We fine-tune the ERes2NetV2 model (Chen et al. 2024a) to obtain improved audio embeddings for our scenario. The training data is based on the MovieStory101 training set, from which we construct a series of pairwise samples: positive pairs consist of two audio segments from the same person, while negative pairs consist of segments from different individuals. In total, we collect 246,834 data pairs, with 5% randomly sampled as the development set, and the rest used for training. The ERes2NetV2 model is fine-tuned using CosineEmbeddingLoss as the loss function. We set the learning rate to 0.005 and the batch size to 32.
@@
Additionally, our audio embedding model outperforms all open-source models (CAM++(Wang et al. 2023), Ecapa-tdnn(Desplanques, Thienpondt, and Demuynck 2020) and ERes2NetV2(Chen et al. 2024a)) under this condition, as Table 9 shown.
@@
CAM++ 0.22 0.36
Ecapa-tdnn 0.23 0.36
ERes2NetV2 0.25 0.40
@@ Appendix E Training Details
Pre-training the audio module. In the pre-training stage, we use the same pre-training dataset of SALMONN (Tang et al. 2023), consisting of both 960-hour LibriSpeech training set (Panayotov et al. 2015) and 1000-hour GigaSpeech M-set (Chen et al. 2021) for speech recognition, as well as 2800-hour WavCaps (Mei et al. 2024), AudioCaps (Kim et al. 2019) and Clotho (Drossos, Lipping, and Virtanen 2020) for audio captioning. In this phase, we only pre-train the window-level Q-former, while freezing the audio encoders and LLM. For the window-level Q-former, we use only one trainable query, and set $L=17$, which corresponds to approximately 0.33 seconds per window. This configuration outputs 88 textual tokens from the Q-former for a 30-second audio segment. The model was trained on 16 H800 GPUs for 50,000 steps using a batch size of 6.
Fine-tuning on audio and visual tasks. To enhance the model’s ability in audio and visual character identification, we collect training data for the following tasks: audio diarization, and character recognition, identification, and tracking. For audio diarization, we sampled data from AMI (Mccowan et al. 2005), AISHELL-4 (Fu et al. 2021), AliMeeting (Yu et al. 2022a, b), CallHome (Przybocki and Martin 2001), DIHARD (Ryant et al. 2019), VoxConverse (Chung et al. 2020), and LibriCSS (Chen et al. 2020), creating a training dataset consisting of 290,088 minute-level utterances. Additionally, using MovieNet (Huang et al. 2020b), we construct training data for three tasks: (1) character recognition, which involves recognizing characters from frames; (2) character identification, where given a cast list, the model identifies the characters in the frames; and (3) character tracking, where the model tracks a character’s position across all frames based on a given face. We collect 33,576 data for these three tasks.